OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <3, 0>, v = <0, -6> Orthogonal Neither Parallel

OpenStudy (unklerhaukus):

what do you get when you take the dot product?

OpenStudy (unklerhaukus):

the dot product of orthogonal vectors is always zero

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got Neither

OpenStudy (phi):

do you know how to do a dot product ?

OpenStudy (phi):

multiply "corresponding" numbers and add them up

OpenStudy (anonymous):

the dot product is given by : u*v = u1v1 + u2v2

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

sqrt(0^2 + -6^2) = 6i sqrt(3^2 + 0^2) =3

OpenStudy (phi):

that looks like you tried to find the length of each vector (and by the way -6*-6 is +36 so you get 6 not 6i ) but you want to do a dot product. multiply the first # times the first # and 2nd # times the 2nd # then add

OpenStudy (anonymous):

3*0 + 0*-6 =0

OpenStudy (unklerhaukus):

the magnitude (or lenght) is when you take the dot product of a vector with it self, which looks like what youve tied to do there, but for this question you have to take the dot product of u with v \[\vec u\cdot\vec v=\langle u_1,u_2\rangle\cdot\langle v_1,v_2\rangle=u_1v_1+u_2v_2\].... like that

OpenStudy (phi):

yes, and that is *very interesting* 0 means the vectors are orthogonal

OpenStudy (unklerhaukus):

yeah.

OpenStudy (anonymous):

lol ur name (UnkleRhaukus) i love boondocks

OpenStudy (unklerhaukus):

another way to do this problem would be to draw the vectors,